OpenStudy (anonymous):

How much compound interest will $50,000 earn in 10 years at 3.4% APR compounded yearly, providing no deposits or withdrawals are made? Hint: Use the formula A = P(1.00 + r)n.

OpenStudy (anonymous):

I've gotten 2 different answers and none of them are the choices.

OpenStudy (anonymous):

sorry not 3.4 but 0.034

OpenStudy (anonymous):

A = 50000(1 + 0.034)*10

OpenStudy (anonymous):

sorry for that.

OpenStudy (anonymous):

Okay, so now I got 517000.

OpenStudy (anonymous):

But that isn't an answer choice.

OpenStudy (anonymous):

yes i guess it should be ^10 and not *10 A = 50000 * (1 + 0.034)^10

OpenStudy (anonymous):

that should be something like 69851.44

OpenStudy (anonymous):

Alright haha, I got that now. Thank you. :) @Coolsector

OpenStudy (anonymous):

sorry for the confusion though lol

OpenStudy (anonymous):

Haha it's okay

OpenStudy (agent0smith):

A = P(1.00 + r)^n... not multiplied by n.